# Atomic Localization Fluorescent Microscopy

### v1

@@ IV Large-scale DIGIT
Fig. 4a presents a hyperspectral image of widefield PLE, spanning spatial $(x,y)$ and atomic transition frequency $\omega/2\pi$. We applied intensity-based region grouping to the PLE EMCCD image to distinguish individual emitters whose transitions range from $406.69$ to $406.73$ THz. Fig. 4b shows the reconstructed cluster density map, with each cluster color-coded by the emitter counts $M$ identified through PLE spectrum (See supplementary Information Sec. V). A detailed flowchart of the widefield PLE data analysis can be found in Extended Data Fig. 4.
@@
Next we integrated $8000$ frames at each resolved frequency $\omega_{i}$ and measured the localization uncertainty $\sigma$ for each emitter, as summarized in Fig. 4c. Only CCs exhibiting sub-Ångström $\sigma$ were counted for subsequent analysis. Emitters that did not meet this requirement have typically reduced brightness, exhibit blinking behavior [21], or experience non-uniformity widefield illumination [25], all of which contribute to worse localization uncertainty.
Figs. 4d and 4e present both numerical and experimental studies of the localization error enhancement $\delta_{P}/\delta_{f}$ across varying cluster densities. In Fig. 4d, we evaluate DIGIT’s performance with different emitter counts $M$ within a cluster. The simulation predicts that $\delta_{p}/\delta_{f}=1/M$, indicating that the localization error can be effectively compensated for by higher emitter counts. This is because additional emitters help to refine lattice alignment through MLE. Experimentally, we examined three cluster densities $M=3,4,5$, each closely matching the simulation prediction. This result confirms the collective advantage of dense-implanted clusters.
In Fig. 4e, we investigate $\delta_{P}/\delta_{f}$ for clusters with varying spatial separations $L$ at $M=5$. Numerical simulations show that DIGIT’s advantage diminishes as L grows, because distinguishing the maximum likelihood lattice point from adjacent points requires increasingly precise estimates of the rotation angle $\theta$, as $L*\theta\leq a$. Achieving such high precision poses significant challenges for MLE methods due to computational and experimental limitations. Experimentally, this scenario is further complicated by strain-induced variations in lattice coherence length, which can result in uneven lattice structures and gaps that impede precise localization [26].
Nevertheless, by extending DIGIT to $>600$ emitters across multiple clusters, we demonstrate its scalability and efficiency for parallel analysis. In contrast to scanning-based approaches that image each emitter individually, DIGIT processes many emitters simultaneously, achieving high-throughput localization.
V Discussion and Outlook
@@
As the localization precision of conventional super-resolution microscopy approaches the length scale of electronic orbitals coupled to optical fields, it becomes essential to incorporate the atomic structure of samples as a Bayesian prior. Our demonstration of the DIGIT was made possible by achieving localization precision ($\sigma$) comparable to the lattice constant ($a$), allowing emitters to be accurately mapped to discrete lattice points rather than being confined to a continuous spatial domain. Furthermore, DIGIT provides a new precision scaling law: $\delta_{P}\propto e^{-\sqrt{N}}$ that is isotropic in all dimensions. Besides, parallelized DIGIT demonstrates its scalability by simultaneously localizing multiple emitters. To further access denser clusters, improved spectral resolution becomes critical. One promising strategy involves operating PLE without a repump laser so that emitters approach lifetime-limited spontaneous emission [27, 28].
@@
Looking ahead, DIGIT has a direct application in diamond-based quantum information processes. For instance, the ability to achieve sub-Ångström localization enables in-situ tracking of emitters’ diffusion and interaction [29, 30, 31, 32], increases spatial storage density [33], and facilitates the generation and control of large cluster states made of locally coupled spins [34, 35]. Additionally, it is readily transferable to other periodic host materials, aiding to identify solid-state quantum memories. Potential candidates include point-like defects [36, 37, 38]; extended electronic orbitals in semiconductor quantum dots [39, 40], or molecules such as dibenzoterrylene, tetracene, or other complexes [41, 42, 43, 44, 45].
@@
Acknowledgments –We thank M.Mazaheri, Q.Peng for valuable input with the digital twin localization model. We thank C.S. Peng, A. Roy, H. Wang, J. Ren, Y. Liu for helpful discussions. We thank M.Mazaheri and V.Saggio for proofreading the manuscript. We acknowledge excellent technical support at the Massachusetts Institute of Technology (MIT) for MIT.Nano. Y.Yu performed FIB implantation, J. Daley performed evaporation medal growth. This work was supported in part by the National Science Foundation (NSF) Science-Technology Center (STC) Center for Integrated Quantum Materials (CIQM) under Grant No. DMR-1231319, by the NSF Engineering Research Center for Quantum Networks (CQN) awarded under Cooperative Agreement No. 1941583, and by the MITRE Moonshot Program. Y.D. acknowledges support by Mathworks fellowship.
Author Contribution –Y.D. conducted diamond fabrication, SiV implantation, digital twin development, localization analysis and measurements of the whole experiment. K.C.C. assisted with diamond fabrication. M.E.T assisted with building the wide-field SMLM, the dipole radiation model and digital twin verification. Q.G, Y.H, and D.R.E assisted with calibrating the camera photoresponse. Y.D and Q.G developed and implemented the Bayesian framework. Y.D, M.E.T and D.R.E conducted the theory proposal of optically resolved lattice structure. Y.D and D.R.E. conceived and supported the project. Y.D., M.E.T. and D.R.E wrote the manuscript with input from all authors.
Competing interests –The authors declare no competing financial interests.
Data availability –All the data that support the findings of this study are included in the Article and its Supplementary Information. Source data are available via figshare at 10.6084/m9.figshare.28281212
Code availability – Code for DIGIT principle and widefield DIGIT is available at: https://github.com/sophiaOnPoint/DIGIT

### v2

@@ II.4 Large-scale DIGIT
We present a hyperspectral image of widefield PLE, spanning spatial $(x,y)$ and atomic transition frequency $\omega/2\pi$, as shown in Fig. 4a. We applied intensity-based region grouping to the PLE EMCCD image to distinguish individual emitters whose transitions range from $406.69$ to $406.73$ THz. We successfully identify the number of emitters $M$ at each sub-diffraction limit spot, as shown in Fig. 4b. A detailed flowchart of the widefield PLE data analysis can be found in Extended Data Fig. 4.
@@
We then integrated $8000$ frames at each resolved frequency $\omega_{i}$ and the measured localization uncertainty $\sigma$ for each emitter is summarized in Fig. 4c. Only SiVs exhibiting sub-Ångström $\sigma$ were counted for subsequent analysis. Emitters that did not meet this requirement have typically reduced brightness, exhibit blinking behavior [37], or experience non-uniformity widefield illumination [38], all of which contribute to worse localization uncertainty. As shown in Fig. 4c, we achieved sub-Ångström localization in $172$ emitters, with a mean localization precision of $0.05$ nm. Those results represent the first demonstration of massively parallel imaging with sub-Ångström precision, enabled by the unique combination of wide-field imaging and DIGIT’s ability to resolve clustered emitters onto discrete lattice sites, thereby enhancing localization probability and precision with large-scale emitters.
Large-scale DIGIT enables the exploration of localization dependence across different clusters in parallel. We first analyze the localization enhancement, defined as $\sigma_{p}/\sigma_{f}$, at various clusters with differing emitter densities. The presence of additional emitters provides improved lattice alignment information through maximum likelihood estimation, resulting in a localization enhancement that scales as $\sigma_{p}/\sigma_{f}\propto 1/M$, as shown in Fig. 4d. We validate this prediction using experimental data. Specifically, we examined three cluster densities, $M=3,4,$ and $5$. Compared to the previous case in Fig. 3, increasing $M$ beyond 3 yields a clear improvement in localization. However, no significant difference was observed between the cases with $M=4$ and $M=5$, which may be attributed to the limited number of clusters with $M=5$ under our FIB conditions and potential lattice damage during the emitter generation process.
The spatial separation $L$ is another key factor in the localization process. When emitters within a cluster are further apart, the localization enhancement provided by DIGIT diminishes. This reduction in performance arises because lattice fitting errors become dominant when distinguishing the maximum likelihood lattice point from adjacent points. In such cases, increasingly precise estimates of the rotation angle $\theta$ are required to satisfy the condition $L\theta\leq a$. We examine the ratio $\sigma_{p}/\sigma_{f}$ for clusters with varying spatial separations at $M=5$. As shown in the top panel of Fig. 4e, there exists a critical separation $L=L_{0}$ at which DIGIT can provide enhancement. For the current parameter set, we theoretically identify $L_{0}=3000a$. Experimental results for $L\geq 3000a$ and $L\leq 50a$ are presented in the bottom panel of Fig. 4e. The average value of $\sigma_{p}/\sigma_{f}$ clearly separates around 1 under these two specific conditions, aligning well with theoretical predictions. Moreover, in experimental settings, this effect is further complicated by strain-induced variations in the lattice coherence length, which can result in uneven lattice structures and discontinuities that hinder precise localization [39].
III Discussion and Outlook
@@
As the localization precision of conventional super-resolution microscopy approaches the length scale of electronic orbitals coupled to optical fields, it becomes essential to incorporate the atomic structure of samples as a Bayesian prior. Our demonstration of the DIGIT was made possible by achieving localization precision ($\sigma$) comparable to the lattice constant ($a$), allowing emitters to be accurately mapped to discrete lattice points rather than being confined to a continuous spatial domain. Furthermore, DIGIT provides a new precision scaling law: $\sigma_{p}\propto e^{-\sqrt{N}}$ that is isotropic in all dimensions. Besides, parallelized DIGIT demonstrates its scalability by simultaneously localizing multiple emitters. To further access denser clusters, improved spectral resolution becomes critical. One promising strategy involves operating PLE without a repump laser so that emitters approach lifetime-limited spontaneous emission [40, 41].
@@
Looking ahead, DIGIT has a direct application in diamond-based quantum information processes. For instance, the ability to achieve sub-Ångström localization enables in-situ tracking of emitters’ diffusion and interaction [42, 43, 25, 44], increases spatial storage density [45], and facilitates the generation and control of large cluster states made of locally coupled spins [46, 47]. Additionally, it is readily transferable to other periodic host materials, aiding to identify molecular-scale configurations. Potential candidates include point-like defects [48, 49, 50]; extended electronic orbitals in semiconductor quantum dots [51, 52], or molecules such as dibenzoterrylene, tetracene [53, 54, 55, 56, 57] and nuclear pore complex [16, 58].
@@
Acknowledgment
We thank M.Mazaheri, Q.Peng for valuable input with the digital twin localization model. We thank C.S. Peng, A. Roy, J. Ren, Y. Liu for helpful discussions. We thank M.Mazaheri, V.Saggio and X. Chen for proofreading the manuscript. We acknowledge excellent technical support at the Massachusetts Institute of Technology (MIT) for MIT.Nano. Y.Yu performed FIB implantation, J. Daley performed evaporation metal growth. This work was supported in part by the National Science Foundation (NSF) Science-Technology Center (STC) Center for Integrated Quantum Materials (CIQM) under Grant No. DMR-1231319, by the NSF Engineering Research Center for Quantum Networks (CQN) awarded under Cooperative Agreement No. 1941583, and by the MITRE Moonshot Program. Y.D. acknowledges support by MathWorks fellowship.
Author Contributions
Y.D. conducted diamond fabrication, SiV implantation, digital twin development, localization analysis and measurements of the whole experiment. K.C.C. assisted with diamond fabrication. M.E.T assisted with building the wide-field SMLM, the dipole radiation model and digital twin verification. Q.G, H.W, Y.H, and D.R.E assisted with calibrating the camera photoresponse. Y.D and Q.G developed and implemented the Bayesian framework. Y.D, M.E.T and D.R.E conducted the theory proposal of optically resolved lattice structure. Y.D and D.R.E. conceived and supported the project. Y.D., M.E.T. and D.R.E wrote the manuscript with input from all authors.
Data availability
All the data that support the findings of this study are included in the Article and its Supplementary Information. Source data are available via figshare at 10.6084/m9.figshare.28281212
Code availability
Code for DIGIT principle and widefield DIGIT is available at: https://github.com/sophiaOnPoint/DIGIT